How many liters of fluorine gas are needed to form 379 L of sulfur hexafluoride gas if the following reaction takes place at 2.00 atm and 273.15 K:
S(s) + 3F₂(g) → SF₆ (g)?
1
Expert's answer
2019-12-09T08:05:41-0500

p = 2 atm = 202650 Pa = 202650 J/m3

V(SF6) = 379 l = 0,379 m3

R = 8,314 J/(mol•K)

T = 273,15 K

__________________________________________

pV=nRT

n(SF6) = PV/RT = 202650•0,379/(8,314•273,15) = 33,82 mol

n(F2)=3•n(SF6) = 3•33,82 = 101,46 mol

V(F2)= nRT/p = 101,46•8,314•273,15/202650 = 1,137 m3 = 1137 liters

Have a nice day! :)